# UINav: A maker of UI automation agents

### v1

@@ 4.2 Referee model
In the action space there is no explicit “done” action. This means that a UINav agent does not stop on its own but instead relies on the environment to terminate a task under appropriate conditions. This is common practice in reinforcement learning. Instead of building task-specific termination logic in the environment, we train a referee model to predict whether a task is completed at each step and what its outcome is. The referee model is trained using the exactly same set of demonstrations as the agent, hence it does not incur extra efforts in data collection.
@@
However, the referee also serves a second purpose. A well-known challenge in demonstration-based systems is that they may require excessive developer time to provide a sufficient number of demonstrations Lau (2009) or that it may be difficult to provide samples that are sufficiently different from each other Myers and McDaniel (2001); Lee et al. (2017). By continuously evaluating the execution of a currently-trained agent and identifying failing tasks, the referee guides users towards collecting new demonstrations only for critical scenarios, thus reducing the number of manual demonstrations.
Figure 4: The architecture of the referee model.
The referee is a recurrent neural network (RNN)-based model ( Figure 4). The attention over Transformer-encoded UI elements is similar to that of the agent model, except that the query is the input utterance concatenated with the action history (the action performed in the previous step and its outcome). Although action history could be derived from previous screen representations, feeding it as input directly makes it less challenging as the referee does not have to learn it. The output of the attention module is then fed into a gated recurrent unit (GRU) Cho et al. (2014). The GRU takes this along with the previous internal hidden state as inputs to predict the current status of the step: (1) SUCCESSFUL: the task is completed and it is successful; (2) FAILED: the task has failed or has reached the maximum number of allowed steps; (3) PENDING: the task is ongoing; or (4) INFEASIBLE: the task cannot be executed (e.g., the task may not be well defined). Failed executions are saved along with all their parameters and passed to the demonstrator (see Figure 2).
4.3 Task workflows
@@
UINav’s focus is on generalizing to different execution environments without requiring an excessive number of demonstrations. However, another large source of variability is the input instruction provided in natural language. To address this problem, as in prior work Liu et al. (2018), we design UINav agents to learn general task workflows rather than specific task instances. We do so by pre-processing utterances to identify sub-string that represent the variables of a task. For example, in Search for tiktok in Google, tiktok is the phrase to search for and can be replaced by other keywords. The variable sub-strings are masked and replaced by placeholders before being encoded, so that the utterance embedding is independent on the specific instances. As a result, there is no need to train with different utterances covering the distribution of variables.
@@
For any utterance, all the replaced sub-strings are included in the list of entities associated with the task. A matching vector is computed for each UI element on the screen and is included in the element attributes passed as input to the agent (see §4.1.1). In the matching vector, each scalar is in the range of $[0,1]$ and computed as the cosine similarity between the text label of the UI element and the corresponding entity string.
@@ 5 Increasing robustness and efficiency
So far we have illustrated how UINav balances automation accuracy and number of required demonstrations. Now we describe the techniques that increase the robustness of UINav agents in the face of system delays, changes in the UI, and variations in task descriptions.
5.1 Action validation and macro actions
@@
Manipulating UI elements of an actual device involves dealing with various system-specific issues. There are unavoidable delays between the time a state is collected from a device and when an action predicted by the agent is ready to be performed. Screens can also be slow at loading or updating, hence an agent needs to wait for them to stabilize. These delays are particularly noticeable on a mobile device. To deal with these issues, rather than modeling this variability through more demonstrations, UINav takes various programmatic measures .
5.1.1 Action validation
@@
Before executing a predicted action, UINav checks whether a referenced UI element is still on the current screen and if so, whether it has changed. In this way, UINav can detect whether the predicted action is still applicable, and, in case of an obsolete action, requests a new prediction.
5.1.2 Macro actions
In UINav, every action is executed as a small program that is composed of lower level operations with status checks. Such a program is referred to as macro. Macro actions abstract the system-specific details, thus making it possible to build cross-platform agents and simplifying the agent’s logic. Each macro action is implemented following a state transition graph. Figure 5 shows the state transition graph for most macro actions that result in screen changes, such as click and back. It starts at S0, and transitions among the other states according to incoming events, such as Action dispatched and Screen changed, and exits either successfully (S6) or with a failure (S5). The graphs of other macro actions are similar.
Another advantage of using macro actions is that they package highly dependent, low-level actions. Figure 6 shows an example. The focus_and_type action (inspired from MiniWoB Shi et al. (2017)) consists of 4 low-level actions: clicking the input field to obtain focus, waiting for the blinking cursor to appear, typing the text in the field, and (optionally) pressing Enter. (Note that large arrows in purple are drawn to highlight interesting areas.)
5.2 Demonstration augmentation
To make UINav resilient to UI changes without increasing the number of demonstrations, UINav automatically augments previous demonstrations by randomizing the attributes of randomly-selected, non-critical UI elements. This teaches the agent that these elements may be safely ignored when selecting a suitable action. A UI element is considered critical for a task if it matches one of the following criteria: (1) It is directly manipulated in a demonstration trajectory, (2) Its text label is mentioned in the utterance, (3) Its text label is one of the common phrases appearing on important elements of a UI navigation flow, e.g., $Menu$, $Home$, $Back$, etc., and (4) Its element is of a specific type, such as $Switch$, $Editable$, $Clickable$, etc.
@@ 5.2 Demonstration augmentation
Non-critical UI elements have their attributes modified with a predefined probability by either (i) replacing the embedding of their text labels with random vectors, or (ii) by adding independent random offsets to the four scalars of their bounding boxes, which is equivalent to randomizing both the position and size of UI elements.
6 UINav Console
To collect demonstrations, we have developed a dedicated application, the UINav Console, that can be seen in the right-half side of the screenshots in Figures 1012 in the Appendix. At each step of a demonstration, a user specifies a macro action, including action type, referenced element, and action argument (if any), and then requests execution of the action.
7 Implementation
We built UINav for the Android platform. However, our design is applicable to other platforms and some of our techniques (e.g., macro-actions and screen representation) are specifically designed to be platform agnostic.
Both the agent and the referee models are implemented in TensorFlow. We employ two inference modes, off-device and on-device. During development we use the Python API of TensorFlow to test the models off-device. Once stable, the models are converted to TensorFlow Lite (tflite) for on-device inference. The agent model has about 320k parameters and its tflite version occupies 1.3MB. The referee model has about 430k parameters and it is 1.8MB large. Both models utilize the same pre-trained language model to encode utterances and texts appearing on screens. We choose the smallest model, L-2_H-128_A-2, of SmallBERT Turc et al. (2019), and convert it to a 17.6MB tflite model. Note that no quantization is applied during the tflite conversion of any of the above models. For efficiency, the sentence encoding computation of the agent and referee models are shared.
Table 2 reports the average execution time per task in milliseconds of the agent, referee, and pre-trained BERT models on a high-end and a low-end Android phone. Note that none of the models utilizes any accelerators. Both of the agent and the referee take just a couple of milliseconds on a high-end device and around 5 milliseconds on a low-end device, while the sentence encoding by BERT dominates the total inference time. The time spent in the agent and the referee is stable on different screens, while the sentence encoding by BERT is roughly linear to the number of strings on a screen plus one (for the utterance). We estimate that most of the sentence encoding by BERT can be avoided as there are many texts repeatedly appearing on different screens, and all the variables are masked before encoding any utterance.
Table 2: Average inference times (msec) of the agent, referee, and pre-trained BERT models.
Device Agent Referee BERT Total
High-end phone 1.98 2.21 262.79 267.00
Low-end phone 4.40 5.24 427.63 437.27
@@ 8 System evaluation
We evaluate UINav with the following goals in mind: (i) quantifying the number of demonstrations required to achieve high success rates in a variety of tasks, (ii) quantifying the benefits of various techniques we have proposed in comparison to baseline systems, and (iii) evaluating the trade-offs between a multi-task and single-task agent.
Figure 7: Number of demonstrations in the training set grouped by tasks.
Figure 7: Number of demonstrations in the training set grouped by tasks.
8.1 Data collection
To train and evaluate UINav we collect a dataset of 4257 human demonstrations across 43 different tasks and 100+ Android apps and websites. Apps and websites are selected based on popularity (i.e., top apps in the Google Play Store and rank of most visited websites, with the exclusion of apps that do not run on emulators). [^1] Some examples are Amazon, Walmart, Google Messages, Gmail, Contacts, Airbnb, CNN, Zillow, wikipedia.org, linkedin.com, craiglist.org, homedepot.com, target.com, etc. Tasks are decided by enumerating common functionality these apps and websites support. Each task is identified by an utterance pattern. The names of the tasks can be seen along the X axis in Figure 7. Demonstrators collect demonstrations for evaluation independently from those for training. For training, we leverage the UINav error-driven approach to demonstrations, while for evaluation we encourage demonstrators to cover as many task variations as possible. It is not uncommon that there are more evaluation episodes than training episodes for a task. Across all tasks, 3661 episodes are used for training and 596 for testing.
During both demonstration collection and online evaluation, our learning environment performs a few random operations at the beginning of each task, including randomly changing pixel densities, font scales, device orientation, and issuing a sequence of random number of clicks on randomly selected UI elements. The purpose is to start a task from a random start state and to diversify data coverage.
It is unlikely that the agent sees a training sample that is identical to one in the evaluation set due to: 1. the random initialization of a task execution environment; 2. the dynamic characteristics of a live system; and 3) the randomization of feature vectors applied to the UI elements of training samples during demonstration augmentation.
Unless explicitly specified, all the statistics in this paper are for a single pair of agent and referee models trained to complete all the tasks.
8.2 Settings and metrics
We perform two types of evaluation: offline and online. The offline evaluation automatically compares the output of the neural networks with those in an evaluation dataset. It records two metrics: (i) step accuracy: the percentage of task steps where the neural networks and the dataset have matching outputs; and (ii) task accuracy: the percentage of tasks with fully matching steps. The online evaluation interacts with a real phone and checks whether the evaluation task completed successfully. Online evaluation is more precise as there could be more than one way to accomplish a task, and making a mistake in a step may not necessarily result in a failure if the agent recover afterwards. On the other hand, online evaluation of an agent requires human involvement, if the referee is not sufficiently accurate. The results of online evaluation can also vary across sessions due to the dynamic nature of a live system. Hence, we employ the following protocol for online evaluation: each evaluation is composed of 5 sessions, with each session containing 100 randomly sampled task descriptions. We compute the average success rate of each session.
8.3 Accuracy of agent and referee models
Through error-driven demonstration collection, developers can train agents with near perfect success rates (i.e., >99%). Figure 7 shows the number of training samples collected to reach such performance. The SEARCH task required over 1700 task demonstrations because it must work for 100 or more different apps and websites. All other tasks are specific to a single app and thus require fewer samples, 30 on average.
Table 3: Accuracy of agent and referee models (obtained using offline evaluation).
Agent Referee
step acc task acc acc
95.8% 90.6% 99.5%
Table 3 reports the average accuracy of the agent and referee models obtained through offline evaluation, when using all the demonstrations in the training set. The agent’s task accuracy is over 90% and the step accuracy above 95%. The referee model generally exhibits higher accuracy than the agent because its task is simpler (predicting 1 out of 4 labels) and its evaluation is not affected by the ambiguity of multiple potentially correct paths.
8.4 Baseline comparison and ablation analysis
In order to compare with existing UI automation baselines, we train and evaluate UINav agents on the publicly available MoTIF dataset Burns et al. (2022). The dataset includes two test splits: (i) app seen task unseen which tests whether a model can generalize to new tasks, and (ii) app unseen task seen which tests whether a model can generalize to new apps. As in the original MoTIF evaluation Burns et al. (2022), we train UINav using the low-level MoTIF instructions as input utterances, and compare against the same baselines: Seq2Seq Shridhar et al. (2019), MOCA Singh et al. (2020), and Seq2Act Li et al. (2020a).
Table 4: Task and step accuracy on the MoTIF dataset. UINav +aug represents UINav with demonstration augmentation (§5.2) while UINav +aug+utt represents UINav +aug with utterance masking (§4.3).
@@ 8.4 Baseline comparison and ablation analysis
Table 4 reports the evaluation results.222‘Task accuracy’ and ‘Step accuracy’ are synonyms of ‘Complete sequence’ and ‘Partial sequence’, respectively, which were used in the MoTIF Burns et al. (2022) and Seq2Act Li et al. (2020a) evaluations. For ablation purposes, we consider three variants of UINav, depending on whether demonstration augmentation (UINav +aug) and utterance masking (UINav +aug+utt) are enabled. The basic UINav model overperforms all three baselines, in all except one case (step accuracy in app unseen and task seen). With demo augmentation (UINav +aug) surpasses all baselines by 7 and 11 percentage points in task accuracy and 8.5 and 0.7 points in step accuracy. This demonstrates the effectiveness of demonstration augmentation in exposing the model to a variety of training conditions which helps it to generalize to new apps and new tasks. Generalizing to new apps is generally harder than generalizing to new tasks, but with the addition of utterance matching UINav achieves 59.6% in task accuracy and 81.9% in step accuracy, well above all baselines.
To evaluate the referee model we use again the MoTIF dataset. The dataset contains task instructions labeled as “feasible” or “infeasible”. Infeasible tasks are those whose step instructions cannot be completed in the given app (e.g., the instruction requires a functionality which is not supported by the app). We compare against the system the authors of the MoTIF dataset propose, which is specifically designed to predict the feasibility or infeasibility of a task. Figure 8 reports the confusion matrix and F1 scores of feasibility/infeasibility predictions. For MoTIF we report results for their best system configuration. [^3] Note that the F1 score reported in their original paper Burns et al. (2022) is incorrect (61.1% instead of 63.5%). The UINav referee predicts 4 states. We map SUCCESSFUL and PENDING to “feasible” and FAILED and INFEASIBLE to “infeasible”. Even though our referee model outputs multiple value states, it produces a significantly better F1 score, 80.9% vs. 63.5%. UINav is especially better in identifying infeasible predictions.
@@
Figure 8: Referee model compared to the MoTIF system Burns et al. (2022) using the MoTIF dataset. For MoTIF we report results for their “best combination”.
@@
By default, UINav agents are multi task, because for on-device applications, it is desirable to keep the total size of all deployed models as small as possible. Our next experiment confirms that this choice is also preferable when the number of training samples is small. We compare the performance of a multi-task agent to that of a single-task agent. We select from our pool the 10 tasks with the largest number of demonstrations. We train one multi-task agent using demonstrations across all 10 tasks and 10 single-task agents using demonstrations from individual tasks. We repeat the training for an increasing number of demonstrations.
As Figure 9 shows, the UINav multi-task agent reaches more than 50% accuracy even with just one demonstration per task. The average accuracy for both multi-task and single-task agents surpasses 80% with 40 demonstrations per task. When the number of demonstrations per task is small, the multi-task agent has a clear advantage, which demonstrates transfer learning across tasks is happening (in line with previous results on MiniWoB Humphreys et al. (2022)). As the number of demonstration per task increases, the accuracy of multi-task agents becomes similar to that of single-task agents, and is slightly higher with over 100 demonstrations per task, probably because the single-task agents suffer less from inter-task interference.
Figure 9: Comparison between multi-task and single-task agents for an increasing number of demonstrations per task.
Figure 9: Comparison between multi-task and single-task agents for an increasing number of demonstrations per task.
9 Limitations

### v2

@@ 4.2 Referee model
In the agent’s action space there is no “done” action. This means that the agent does not stop on its own but instead relies on the environment to terminate a task. This is common practice in reinforcement learning. Instead of building task-specific termination logic, we train a referee model to predict whether a task is completed at each step and what its outcome is. The referee is trained using the exactly same set of demonstrations as the agent, hence it does not incur extra effort in data collection. However, it also serves a second purpose.
@@
A well-known challenge in demonstration-based systems is that they can require excessive developer time to collect a sufficient number of demonstrations Lau (2009) and that it may be difficult to provide samples that are sufficiently different from each other Myers and McDaniel (2001); Lee et al. (2017). By automatically evaluating the execution of a currently-trained agent and identifying failing tasks, the referee guides users towards collecting new demonstrations only for critical scenarios. Failed executions are saved along with all their parameters and passed to the demonstrator.
The neural architecture of the referee model is similar to that of the agent except that it is wrapped in a recurrent neural network to consider the history of actions (see §A.3 for more details). The referee predicts one out of 4 labels: (1) SUCCESSFUL: the task is completed successfully; (2) FAILED: the task has failed or has reached the maximum number of allowed steps; (3) PENDING: the task is ongoing; or (4) INFEASIBLE: the task cannot be executed.
4.3 Utterance masking
@@
UINav’s focus is on generalizing to different execution environments without requiring an excessive number of demonstrations. However, another large source of variability is the input instruction provided in natural language. To address this problem, we design UINav agents to learn general task workflows rather than specific task instances. We do so by pre-processing utterances to identify sub-string that represent the variables of a task. For example, in Search for tiktok in Google, tiktok is the phrase to search for and can be replaced by other keywords. The variable sub-strings are masked and replaced by placeholders before being encoded, so that the utterance embedding is independent on the specific instances. As a result, there is no need to train with different utterances covering the distribution of variables.
@@
For any utterance, all the replaced sub-strings are included in the list of entities associated with the task. A matching vector is computed for each UI element on the screen and is included in the element attributes passed as input to the agent. In the matching vector, each scalar is in the range of [0,1]01[0,1][ 0 , 1 ] and computed as the cosine similarity between the text label of the UI element and the corresponding entity string.
@@ 5 Increasing robustness and efficiency
We have described how UINav helps developers balance accuracy and number of demonstrations. Next, we describe the techniques that increase the agent robustness in the face of system delays, UI changes, and variations in task descriptions.
Action validation and macro actions.
@@
Controlling UIs of an actual device involves dealing with various system issues. There are unavoidable delays between the time a state is collected from a device and when a predicted action is ready to be performed. Screens can also be slow at loading or updating, hence an agent needs to wait for them to stabilize. These delays are particularly noticeable on a mobile device. To deal with these issues, rather than modeling this variability through more demonstrations, we take various programmatic measures.
@@
First, before executing an action, UINav validates it by checking whether a referenced UI element is still on the current screen and if so, whether it has changed. If the action is not applicable anymore, it requests a new prediction.
Second, every action is executed as a small program that is composed of lower level operations with status checks. Such a program is referred to as macro. Each macro is implemented following a state transition graph and it is atomic so that while a macro is running the agent stays idle and changes to the screen are not visible to it. An example of macro action is focus_and_type which comprises 4 low-level actions: clicking the input field to obtain focus, waiting for the blinking cursor to appear, typing the text in the field, and (optionally) pressing Enter. See §A.4 for more details.
Demonstration augmentation.
@@
To further limit the number of required demonstrations and amplify the learning brought by each one, UINav also augments the collected demonstrations by randomizing the attributes of randomly-selected, non-critical UI elements. This teaches the agent which elements may be safely ignored, and ultimately makes it more tolerant to UI changes. Non-critical UI elements have their attributes modified with a predefined probability by either (i) replacing the embedding of their text labels with random vectors, or (ii) by adding random offsets to the four scalars of their bounding boxes, which is equivalent to randomizing both the element’s position and size. Despite its simplicity, demo augmentation is highly effective at improving UINav’s performance (see Table 2).
Table 1: Inference time (msec) on high/low-end phones. None of the models utilize any accelerators.
Device Agent Referee SmallBERT Total
High-end 1.98 2.21 262.79 267.00
Low-end 4.40 5.24 427.63 437.27
@@ 6 System evaluation
We built UINav for Android. Both the agent and referee are implemented in TensorFlow. The agent model has 320k parameters and its tflite version occupies 1.3MB, while the referee has 430k parameters and it is 1.8MB large. For text encoding we use SmallBERT Turc et al. (2019) and convert it to a 17.6MB tflite model. No quantization is applied during the conversion (More implementation details in §A.5). As shown in Table 1), both the agent and referee take only a couple of milliseconds to execute on a high-end phone (e.g., Pixel6pro) and around 5 milliseconds on a low-end phone (Pixel 3a). BERT dominates the total time.
Table 2: Task and step accuracy on MoTIF.
@@
Figure 3: Referee model compared to the MoTIF system Burns et al. (2022) using the MoTIF dataset.
Figure 4: Number of demonstrations in the training set collected for 43 tasks across 128 apps/websites.
6.1 Agent and referee accuracy
We evaluate UINav on the MoTIF dataset Burns et al. (2022). MoTIF includes two splits: (i) app seen task unseen which tests whether a model can generalize to new tasks, and (ii) app unseen task seen which tests whether a model can generalize to new apps. As in the evaluation of the MoTIF system, we train UINav using low-level instructions, and compare against three baselines: Seq2Seq Shridhar et al. (2019), MOCA Singh et al. (2020), and Seq2Act Li et al. (2020). More training details in §A.7. We measure (i) step accuracy, the percentage of task steps where the model and the dataset have matching outputs, and (ii) task accuracy, the percentage of tasks with all steps matching.
Table 2 reports the results. For ablation purposes, we consider three variants of UINav, depending on whether demonstration augmentation (+aug) and utterance masking (+utt) are enabled. UINav+aug surpasses all baselines by 7 and 11 percentage points in task accuracy and 8.5 and 0.7 points in step accuracy. Without demo augmentation UINav outperforms all three baselines, in all except one case (step accuracy in app unseen and task seen). This demonstrates the effectiveness of the UINav design and how demo augmentation effectively exposes the model to a larger variety of training conditions thus improving generalizability. Generalizing to new apps is generally harder than generalizing to new tasks, but with the addition of utterance matching UINav achieves 59.6% in task accuracy and 81.9% in step accuracy, well above all baselines.
To evaluate the referee model we use again the MoTIF dataset as its traces are labeled as “feasible” or “infeasible”, depending on whether the task was successfully completed. We compare against the MoTIF system, specifically designed to predict task feasibility/infeasibility. As the UINav referee predicts 4 states, we map SUCCESSFUL/ PENDING to “feasible” and FAILED/INFEASIBLE to “infeasible”. As Fig. 3 shows, our referee model produces a significantly better F1 score, 80.9% vs. 63.5%, and it is especially better in identifying infeasible tasks.
6.2 Demonstration effort
To evaluate the effectiveness of the error-driven demo collection approach of UINav we cannot use static datasets. Hence, we quantify the demonstration effort of UINav by using it to train high-accuracy agents for 43 different tasks across 128 Android apps and websites, selected based on popularity (e.g., Gmail, Contacts, Amazon, Airbnb, linkedin.com, target.com, etc.). Please see §A.9 for the full list. For demo collection we build a dedicated GUI which can be connected to Android phones or emulators (see §A.6). The GUI supports macro actions and error-driven data collection. During data collection and testing, the environment automatically performs a few random operations at the beginning of each task, including randomly changing pixel densities, font scales, device orientation, and issuing a sequence of random number of clicks on randomly selected UI elements. The purpose is to start a task from a random state and to diversify data coverage.
We collect demonstrations with the goal to achieve near perfect success rates. With the exception of the search task we collect from 10 to 106 demonstrations (on average 32.7) per task, 3661 in total (Fig. 4). Collecting 10 demonstrations takes less than 10 minutes. The search task must work across 100+ apps hence requiring 1700+ samples. To verify this data is sufficient to train accurate agents, in a second phase we collect additional 596 test samples. Because of the random initialization of the environment, and the dynamic characteristics of a live system, it is unlikely that the models see a training sample that is identical to a test one. The UINav agent achieves 90.6% task accuracy and 95.8% step accuracy; the referee is 99.5% accurate.
@@ 6.2 Demonstration effort
To reduce the resource overhead on mobile devices, we train a single multi-task agent. We show this choice is preferable also for small numbers of demonstrations. From our in-house dataset, we select the 10 tasks with the largest number of demonstrations. We then train one multi-task UINav agent using demonstrations across all 10 tasks and 10 single-task UINav agents using demonstrations from individual tasks. We repeat the training for an increasing number of demonstrations. As Fig. 5 shows, the multi-task agent reaches 51% accuracy even with just one demonstration, demonstrating transfer learning across tasks is happening. The average accuracy for both multi-task and single-task agents surpasses 80% with 40 demonstrations.
Figure 5: Comparison between multi- and single-task agents with an increasing number of demonstrations.
7 Conclusions
We presented a demonstration-based system for building small and fast UI automation agents that are suitable for mobile devices. Our approach requires small human effort and no coding skills. With modest numbers of demonstrations UINav agents achieve near perfect success rate on previously-seen tasks and with more effort they can generalize well to new tasks and applications.
Ethical considerations
A use case that motivates UINav agents include screen readers for visually-impaired users. As accessibility labels are often missing or incomplete in mobile apps, UINav can provide them with access to a much wider range of applications and functionality. Another potential use case of UINav is task automation, which has societal, security and privacy implications. An agent may leak private information or carry out a task in an unacceptable way or produce unwanted side effects. Malicious actors could also use UINav agents for undesired purposes such as overriding anti-fraud mechanisms or manipulating applications to achieve undesirable goals.
To develop UINav we collected a dataset internally. The demonstrators were asked to avoid entering any private information and received fair compensation.
Appendix A Appendix
Limitations